pls. help me solve this prob. tnx.
prob.1. A car and a lorry set off on a 390-km journey at the same time. The lorry traveled at a speed of 60km/hr. The car reached the destination 1 and 1/2 hour earlier than the lorry. find the speed of the car.

Answer by stanbon(75887)

A car and a lorry set off on a 390-km journey at the same time. The lorry traveled at a speed of 60km/hr. The car reached the destination 1 and 1/2 hour earlier than the lorry. find the speed of the car.
-----
Lorry DATA:
rate = 60 km/hr ; distance = 390 km ; time = d/r = 390/60 = 13/2 hrs
----
Car DATA:
distance = 390 km ; time = (13/2)-(3/2) = 5 hrs
---
Car rate = distance/time = 390/5 = 78 km/hr
=================================
Cheers,
Stan H.
===========